In Wilburts math class his garde is based on the average of six tests,each wirth 100 points.Wilburt always worries about his grades.He knows that his average on the firsr four tests is 88.5. What is the lowest average he could get on his next two tests abd still average 90 overall?
Answer by stanbon(75887)

In Wilburts math class his grade is based on the average of six tests, each with 100 points.
Wilburt always worries about his grades. He knows that his average on the firsr four tests is 88.5.
Total points on the first four tests: 4*88.5 = 354
--------------------------------------------------------------
What is the lowest average he could get on his next two tests abd still average 90 overall?
----
To average 90 on 6 tests he would need 6*90 = 540 points.
----
The last two must total 540-354 = 186 points
------
Since there are 2 tests they must average 186/2 = 93 points.
-----------------------------------------
Cheers,
Stan H.
===================================